# Diversity as a Reward: Fine-Tuning LLMs on a Mixture of Domain-Undetermined Data

### v1

@@ B.8 Prompts for the Generation Stage
Appendix C Proof of Section 4
C.1 Proof of Proposition 4.3
Lemma C.1 (Cross-Domain Risk Decomposition).
Under Assumption 4.1, the average cross-domain risk becomes:
$\mathbb{E}_{k\sim\mathcal{K}}\mathcal{L}_{\mathcal{D}_{k}}\left(\textstyle\sum_{m=1}^{M}\pi_{km}h_{\theta_{m}}\right)=\sum_{m=1}^{M}\lambda_{m}\mathbb{E}_{\tilde{\mathcal{D}}_{m}}\left[l\left(\textstyle\sum_{m^{\prime}=1}^{M}\pi_{km^{\prime}}h_{\theta_{m^{\prime}}}(x),y\right)\right],$ (21)
where $\lambda_{m}=\mathbb{E}_{k}[\pi_{km}]$ represents global activation of capability $m$.
Proof.
Under the finite variance condition $\mathbb{E}_{\tilde{\mathcal{D}}_{m}}[\|l\|_{2}^{2}]<\infty$, Fubini’s theorem holds, and results in:
$\mathbb{E}_{k}\mathcal{L}_{\mathcal{D}_{k}}$ $=\mathbb{E}_{k}\mathbb{E}_{(x,y)\sim\mathcal{D}_{k}}\left[l\left(\textstyle\sum_{m}\pi_{km}h_{\theta_{m}}(x),y\right)\right]$
$=\mathbb{E}_{k}\mathbb{E}_{m\sim\pi_{k}}\mathbb{E}_{\tilde{\mathcal{D}}_{m}}\left[l\left(\textstyle\sum_{m^{\prime}}\pi_{km^{\prime}}h_{\theta_{m^{\prime}}}(x),y\right)\right]$
$=\sum_{m=1}^{M}\lambda_{m}\mathbb{E}_{\tilde{\mathcal{D}}_{m}}\left[l\left(\textstyle\sum_{m^{\prime}}\pi_{km^{\prime}}h_{\theta_{m^{\prime}}}(x),y\right)\right]$ (22)
Theorem C.2 (Global Optimality Condition).
Under Assumptions 4.1-4.2, the optimal predictors $\{h_{\theta_{m}^{*}}\}$ minimize the cross-domain risk when:
$\forall m,\ \theta_{m}^{*}=\mathop{\mathrm{argmin}}_{\theta}\mathbb{E}_{\tilde{\mathcal{D}}_{m}}[l(h_{\theta}(x),y)].$ (23)
Proof.
Taking derivatives of Lemma C.1’s right-hand side:
$\frac{\partial}{\partial\theta_{m}}\sum_{m^{\prime}=1}^{M}\lambda_{m^{\prime}}\mathbb{E}_{\tilde{\mathcal{D}}_{m^{\prime}}}[\cdot]=\lambda_{m}\mathbb{E}_{\tilde{\mathcal{D}}_{m}}\left[\frac{\partial l}{\partial h}\frac{\partial h_{\theta_{m}}}{\partial\theta_{m}}\right]+\sum_{m^{\prime}\neq m}\lambda_{m^{\prime}}\mathbb{E}_{\tilde{\mathcal{D}}_{m^{\prime}}}\left[\frac{\partial l}{\partial h}\frac{\partial h_{\theta_{m}}}{\partial\theta_{m}}\right].$
By strong convexity and optimality of $\theta_{m}^{*}$, all terms vanish when $\theta_{m}=\theta_{m}^{*}$. ∎
C.2 Proof of Proposition 4.4
Proof of Inter-Diversity
To prove the inter-diversity decomposition as given in Proposition 4.4, we begin by considering the inter-diversity metric $\Phi_{\text{inter}}$, defined as the expected pairwise distance between domain centroids from Eq (3):
$\Phi_{\text{inter}}=\mathbb{E}_{k\neq l}\left[\|\mathcal{C}_{k}-\mathcal{C}_{l}\|_{2}\right].$ (24)
Substituting the expression for each domain centroid (Eq (10)) $\mathcal{C}_{k}=\sum_{m=1}^{M}\pi_{km}\mathcal{C}_{m}$, we have:
$\|\mathcal{C}_{k}-\mathcal{C}_{l}\|_{2}=\left\|\sum_{m=1}^{M}\pi_{km}\mathcal{C}_{m}-\sum_{n=1}^{M}\pi_{ln}\mathcal{C}_{n}\right\|_{2}.$ (25)
Applying the triangle inequality and expanding, the distance becomes:
$=\left\|\sum_{m=1}^{M}\sum_{n=1}^{M}(\pi_{km}-\pi_{ln})\mathcal{C}_{m}+\pi_{ln}\mathcal{C}_{n}\right\|_{2}.$ (26)
Taking the expectation over all domain pairs $k\neq l$ and using linearity, we have:
$\Phi_{\text{inter}}=\mathbb{E}_{k\neq l}\left[\sum_{m=1}^{M}\sum_{n=1}^{M}(\pi_{km}\pi_{ln})\|\mathcal{C}_{m}-\mathcal{C}_{n}\|_{2}\right].$ (27)
Recognizing that $\lambda_{m}=\mathbb{E}_{k}[\pi_{km}]$ is the expected activation of capability $m$, we can simplify further:
$=\sum_{m=1}^{M}\sum_{n=1}^{M}\lambda_{m}\lambda_{n}\|\mathcal{C}_{m}-\mathcal{C}_{n}\|_{2}.$ (28)
This completes the proof, showing that the inter-diversity metric is determined by the expected pairwise distances between latent capability centroids, weighted by their activation probabilities.
Proof of Intra-Diversity
The intra-diversity metric for domain $k$, denoted as $\Phi_{\text{intra}}^{(k)}$, measures the variance of data samples within the domain relative to its centroid $\mathcal{C}_{k}$. We aim to show that:
$\Phi_{\text{intra}}^{(k)}\leq\sum_{m=1}^{M}\pi_{km}\|\mathcal{C}_{m}-\mathcal{C}_{k}\|_{2}^{2}+\mathbb{E}_{m\sim\pi_{k}}[\text{Var}(\tilde{\mathcal{D}}_{m})].$ (29)
Proof.
To prove this bound, we start by expressing the intra-diversity metric $\Phi_{\text{intra}}^{(k)}$ as the expected squared distance from each sample $\mathbf{x}_{i}^{(k)}$ to the domain centroid $\mathcal{C}_{k}$ from Eq (5):
$\Phi_{\text{intra}}^{(k)}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\|\mathbf{x}_{i}^{(k)}-\mathcal{C}_{k}\|_{2}^{2}.$ (30)
Using the definition of the domain centroid (Eq (10)) $\mathcal{C}_{k}=\sum_{m=1}^{M}\pi_{km}\mathcal{C}_{m}$ in , we can decompose the squared distance for each sample:
$\|\mathbf{x}_{i}^{(k)}-\mathcal{C}_{k}\|_{2}^{2}=\|\mathbf{x}_{i}^{(k)}-\sum_{m=1}^{M}\pi_{km}\mathcal{C}_{m}\|_{2}^{2}.$ (31)
Applying the variance decomposition, we have:
$\|\mathbf{x}_{i}^{(k)}-\mathcal{C}_{k}\|_{2}^{2}=\|\mathbf{x}_{i}^{(k)}-\sum_{m=1}^{M}\pi_{km}\mathcal{C}_{m}\|_{2}^{2}=\sum_{m=1}^{M}\pi_{km}\|\mathbf{x}_{i}^{(k)}-\mathcal{C}_{m}\|_{2}^{2}+\sum_{m=1}^{M}\pi_{km}\|\mathcal{C}_{m}-\mathcal{C}_{k}\|_{2}^{2},$ (32)
where the first term represents the variance within each latent capability, and the second term accounts for the distance between latent capability centroids and the domain centroid.
By taking the expectation over the samples in domain $k$, we get:
$\Phi_{\text{intra}}^{(k)}\leq\sum_{m=1}^{M}\pi_{km}\|\mathcal{C}_{m}-\mathcal{C}_{k}\|_{2}^{2}+\mathbb{E}_{m\sim\pi_{k}}[\text{Var}(\tilde{\mathcal{D}}_{m})],$ (33)
where $\text{Var}(\tilde{\mathcal{D}}_{m})$ represents the variance within the $m$-th latent capability. This completes the proof of the intra-diversity bound. ∎
Appendix D More Experimental Results
D.1 Visualization of Embeddings on Llama3.1-8B & Qwen2.5-7B
We process concatenated data samples with “instruction”+“input”+“output” pairs through each LLM’s embedding layer, computing mean token embeddings for visualization. Beyond Qwen2-7B results in Fig. 1, we present the t-SNE visualization of data samples from Llama3.1-8 and Qwen2.5-7B with different distributions in Fig. 8 and Fig. 9. It is evident that, despite differences in architecture or model, the method of filtering data through Inter-Diversity and Intra-Diversity is effective. Notably, although the embedding distributions of different models are not identical, they exhibit similar behavior, indicating that the representations learned by the embedding layers are comparable.
Figure 8: The t-SNE visualization of embeddings for data samples with different distributions on Llama3.1-8B. (a) The data pool of all 40,000 samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids based on Inter-Diversity, (d) Distribution of data closest to other domain centroids based on Inter-Diversity, (e) Distribution of data closest to its own domain centroid based on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid based on Inter-Diversity.

### v2

@@ C.2 Implementation of Model-Aware Training Data Generation
Appendix D Details of Baselines
D.1 Implementation of Baselines
Random Selection(Rand)
Instruction Length (IL)
The length of the instruction can be considered a measure of input complexity. It is widely believed [5, 53] that more complex data is beneficial for enhancing model capabilities. Therefore, we select a subset of 8,000 entries with the maximum number of words (based on spaces) in the concatenation of Instruction and Input as part of the filtering process.
AlpaGasus [10]
[…]
Instruction: [Instruction]
Input: [Input]
Response: [Response]
User Prompt:
Figure 3: The distribution of the score by Alpagasus.
Instag [30]
[…]
[begin]
{Instruction + Input}
[end]
User Prompt:
Please provide coarse-grained tags, such as "Spelling and Grammar Check" and "Cosplay", to identify main intentions of above instruction. Your answer should be a list including titles of tags and a brief explanation of each tag. Your response have to strictly follow this JSON format: ["tag": str, "explanation": str]. Please response in English.
Figure 4: The distribution of the tags by Instag.
SuperFilter [26]
Deita [29]
D.2 Visualization of Baseline’s Selected Data
We project all baseline-selected data samples through Llama3.1-8B’s, Qwen2-7B’s and Qwen2.5-7B’s embedding layer and visualize their distributions via t-SNE dimensionality reduction in Fig. 5, Fig. 6, and Fig. 7, respectively. This visualization provides intuitive insights into how different data selection strategies handle domain-mixed data. Three key observations and conjectures emerge:
  • Certain methods (particularly Instruction Length, Deita-C, and Deita-Q) demonstrate pronounced distributional skew toward coding-domain samples, which may partially explain their suboptimal HumanEval performance shown in Table 20.
  • The domain-agnostic nature of baseline approaches reveals fundamental limitations in preserving original data distributions when processing domain-unspecified samples, as evidenced by the visualization.
  • Although distributional shifts are evident, the absence of commensurate performance deterioration across all methods empirically validates our theoretical framework (Section 3.4), which establishes connections between data distribution shift.
Figure 5: t-SNE visualization of data samples selected by different baselines using Llama3.1-8B embeddings. While original labels were removed during training, we preserve them for interpretability.
Figure 6: t-SNE visualization of data samples selected by different baselines using Qwen2-7B embeddings. While original labels were removed during training, we preserve them for interpretability.
Figure 7: t-SNE visualization of data samples selected by different baselines using Qwen2.5-7B embeddings. While original labels were removed during training, we preserve them for interpretability.
Appendix E Derivation of Proposition 3.2
We derive the density ratio by systematically expanding and substituting terms:
$\frac{q(\mathbf{x})}{p(\mathbf{x})}=\frac{\sum_{k=1}^{K}q(c=k)q(\mathbf{x}|c=k)}{p(\mathbf{x})}$ (10)
Expand $q(\mathbf{x})$ via the law of total probability over domains.
$=\frac{\sum_{k=1}^{K}q(c=k)p(\mathbf{x}|c=k)}{p(\mathbf{x})}$ (11)
Apply Assumption 3.1 ($q(\mathbf{x}|c=k)=p(\mathbf{x}|c=k)$) to unify domain-conditional densities.
$=\frac{\sum_{k=1}^{K}q(c=k)\frac{p(c=k|\mathbf{x})p(\mathbf{x})}{p(c=k)}}{p(\mathbf{x})}$ (12)
Express $p(\mathbf{x}|c=k)$ via Bayes’ rule inversion: $p(\mathbf{x}|c=k)=\frac{p(c=k|\mathbf{x})p(\mathbf{x})}{p(c=k)}$.
$=\sum_{k=1}^{K}\frac{q(c=k)}{p(c=k)}p(c=k|\mathbf{x})\cdot\frac{p(\mathbf{x})}{p(\mathbf{x})}$ (13)
Factor out $p(\mathbf{x})$ from the numerator and the denominator.
$=\sum_{k=1}^{K}\underbrace{\frac{q(c=k)}{p(c=k)}}_{\lambda_{k}}p(c=k|\mathbf{x})$ (14)
Define domain proportion ratio $\lambda_{k}:=\frac{q(c=k)}{p(c=k)}$.
This demonstrates that the density ratio decomposes linearly into domain-specific components weighted by their proportion shifts ($\lambda_{k}$) and posterior probabilities ($p(c=k|\mathbf{x})$).
Appendix F Comprehensive Experiments
F.1 Performance on Comprehensive Benchmark MMLU
The benchmarks selected in Section 3.1 are specifically tailored to correspond with the domain-specific capabilities of the data sources. To comprehensively evaluate the performance of DaaR across broader assessment metrics and more extensive benchmarking frameworks, we adopted MMLU as an out-of-distribution (OOD) evaluation protocol, detailed description as:
  • MMLU is a benchmark designed to evaluate the multitask capabilities of language models across diverse subjects, covering 57 tasks spanning topics with questions ranging from elementary to professional levels.
Table 6: Evaluation results of Qwen2-7B and Qwen2.5-7B on MMLU.
Models Methods Humanities STEM Social Science Other MMLU-Avg
Qwen2-7B Raw 38.58 25.77 24.94 27.82 28.98
Random 70.02 58.54 80.00 72.26 68.81
Instruction Len 71.34 58.09 79.77 70.71 68.55
Alpagasus [10] 56.55 49.10 71.88 65.54 59.34
Instag-C [30] 70.28 58.69 79.92 71.17 68.65
Instag-D [30] 67.87 58.35 79.82 71.61 68.07
SuperFilter [26] 52.55 49.16 57.99 60.01 54.27
Deita-C [29] 68.05 58.53 80.12 70.97 68.08
Deita-D [29] 67.11 58.22 79.17 72.13 67.83
Deita-Q [29] 56.79 52.82 69.01 63.06 59.47
DaaR (Ours) 70.49 59.94 79.67 72.74 69.42
Qwen2.5-7B Raw 73.41 48.96 72.79 58.46 61.72
Random 75.40 64.47 81.46 72.71 72.42
Instruction Len 75.08 63.74 81.40 72.96 72.15
Alpagasus [10] 75.39 64.85 81.73 73.13 72.70
Instag-C [30] 75.26 64.35 81.64 73.03 72.46
Instag-D [30] 74.83 65.05 81.48 73.16 72.59
SuperFilter [26] 75.24 66.34 82.19 73.99 73.45
Deita-C [29] 75.76 64.18 81.59 72.83 72.46
Deita-D [29] 75.27 65.60 81.55 72.99 72.85
Deita-Q [29] 75.81 64.14 81.41 73.65 72.61
DaaR (Ours) 74.91 65.84 81.85 73.68 73.07
As demonstrated in Table 6, despite not being explicitly optimized for broader capabilities during its design phase, DaaR still demonstrates competitive performance in comprehensive capability assessments. The method achieved the highest score on Qwen2-7B, surpassing the second-place Random selection by 0.61 accuracy points. On Qwen2.5-7B, it exhibited performance closely following SuperFilter, thereby confirming the generalization potential of DaaR to a certain extent.
Notably, Llama3.1-8B was excluded from reference comparisons due to significant result fluctuations observed during Opencompass evaluations (as shown in Table 7 that the results range from 0.04 to 42.52). We further identified inconsistent failures in model prediction generation during the evaluation process, which compromised the reliability of performance measurement. These implementation challenges prevented valid comparative analysis with Llama3.1-8B in our experiments.
Table 7: Invalid evaluation results of Llama3.1-8B on MMLU.
Models Methods Humanities STEM Social Science Other MMLU-Avg
Llama3.1-8B Raw 0.02 0.01 0.08 0.07 0.04
Random 30.13 15.34 30.73 27.51 24.73
Instruction Len 23.38 11.65 20.43 18.26 17.68
Alpagasus [10] 35.99 23.28 36.80 34.37 31.56
Instag-C [30] 25.87 12.58 33.73 23.71 22.60
Instag-D [30] 42.50 19.44 47.70 37.19 34.70
SuperFilter [26] 11.61 9.40 12.11 11.71 11.00
Deita-C [29] 8.79 0.97 0.60 2.27 2.97
Deita-D [29] 50.79 29.76 53.04 43.21 42.52
Deita-Q [29] 33.99 23.36 26.31 24.27 26.61
DaaR (Ours) 45.18 20.94 46.47 34.20 34.87
F.2 Validation on Qwen3-8B
As a prevailing trend in the development of current LLMs, the RL-based (reinforcement learning) thinking model has demonstrated significant capabilities and potential (e.g. DeepSeek series [28], Llama4 series and Qwen3 series). To validate the efficacy of DaaR on the most up-to-date and advanced LLMs, we conducted a series of verification experiments using Qwen3-8B  [^2] https://github.com/QwenLM/Qwen3. It should be noted that, since Qwen3 does not adhere to conventional SFT paradigms, as well as the comprehensive support for training & evaluation of the Qwen3 series remains incomplete, the experiments presented in this section primarily serve as exploratory investigations. A comprehensive analysis of data diversity across various training methodologies and LLM architecture will be reserved for future research endeavors.
Observation Studies on Qwen3
Following the methodology delineated in Section 3, we employed Qwen3’s embedding layer to process the identical data pool. The t-SNE visualization of Qwen3’s embeddings (Fig. 8) demonstrates comparable domain discrimination capabilities.